OpenStudy (anonymous):

TR= 440Q-3Q^2 TC=14Q+225. Find max profit?

OpenStudy (anonymous):

I think you need to find the marginal cost (MC) and marginal revenue (MR) of the two formulas. To maximize profits, MC=MR. You get the marginals by differentiating. SO MR= 440-6q MC= 14 Then 440-6q=14 440-14=6q 426=6q 71=q So then the max output is 71 and then the profit output will be TR-TC So X= 440Q-3Q^2-14Q-225 X= 426q - 3q^2 - 225 X=30246 - 15348 X=14898 So mazimum profit is 14898.